# Working Backwards: Learning to Place by Picking

### v2

@@ III-C2 Placing by reversing
An example of a place trajectory is shown in Figure 1. Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}_{\bar{M}-i-1},$ (4)
where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in SE(3)$ is a relative change in EE pose. We define the expert action as $\mathbf{a}_{i}=(\Delta\mathbf{T}_{i},a_{\text{gripper},i})$, where $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}-1}$, and store the sequences of observations and actions as an expert trajectory. We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$. We downsample the images to a resolution of $128\times 128$ and also stack the three previous RGB frames, which results in $\mathbf{o}_{i}\in\mathbb{R}^{128\times 128\times 12}$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper (i.e. $\Delta\mathbf{T}_{i}=\boldsymbol{\mathbf{I}}$ and $a_{\text{gripper},i}=0$) for N time steps. We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples
@@
III-D Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise. We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [33], and actions [34]. We also draw inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [35], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment. We call this variant noise-augmented data collection. In practice, we represent pose commands $\Delta\mathbf{T}$ as a combination of a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a rotation vector $\boldsymbol{\theta}\in\mathbb{R}^{3}$. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We perturb the translation of the pose with
@@ III-D Noise-Augmented Data Collection
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation with
@@
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We perturb both the rotation axis and the rotation angle. We find $\sigma_{t}=0.5~{}$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5~{}$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-E Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP. We use behavioral cloning (BC) [36] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}\,|\,\mathbf{o}_{t})}.$ (8)
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [37], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [38] and [39], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value of $61e-$ 6. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss.
@@ IV Experiments
[…]
@@
We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 5. In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we present our quantitative results for the more challenging dishrack loading task as it better represents the type of task, with additional contact constraints from the environment, that PvP was designed for. We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
@@ IV-A PvP Data Collection Robustness
We require a robust data collection loop to collect a large number of place demonstrations autonomously with PvP. We studied the effects of compliant control for grasping (CCG) and tactile regrasping (TR), as introduced in Section III, on the robustness of PvP.
We observed two main causes, initially discussed in Section III-A2, that led to misplaced plates during data collection. First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot. With PvP, we mitigate this failure mode with CCG. The second failure mode involved unstable grasps that lead to a large relative motion between the object and the gripper. With PvP, we mitigate unstable grasps using TR to preemptively detect unstable grasps and to regrasp if needed.
We conducted an ablation study on CCG and TR where we ran self-supervised data collection for a total of 128 episodes and recorded the number of failures. We visualize the results in Figure 6. The naive approach without CCG and without TR failed a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The amount of failures dropped to three with the addition of CCG, which raises the average to $43$ episodes collected before each intervention. Finally, with both CCG and TR, we were able to successfully collect 128 episodes without any human intervention.
@@ IV-B Noise Augmentation Ablation
TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation using two different policy action representations. The noise-augmented data collection procedure improves the performance of both deterministic and Gaussian mixture policies.
@@
[…]
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the action allows the policy to better capture this by not having to average over multiple possible expert actions [40]. We also qualitatively observed that policies trained with noise-augmented datasets were better at re-adjusting the object with respect to the goal.
@@ IV-C Comparison to Kinesthetic Teaching
We compared the performance of policies trained using demonstration data from PvP and from kinesthetic teaching. In particular, we investigated whether the quality of the demonstrations differ. To compare, we trained two separate policies, both sharing the architecture outlined in Section III-E, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [41]. We used a Gaussian mixture model as the action representation of both policies. For PvP, we used the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
[…]
We hypothesize that the kinesthetic policy performed more poorly than the PvP policy because the overall quality of the human demonstrations was lower than the programmatic or machine-generated demonstrations produced by PvP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [42, 43, 39]. In particular, human demonstrations can be sub-optimal due to mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think). The authors of [39] note that the relative average trajectory length is a good proxy for the quality of the dataset. The average number of time steps for each demonstration was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset. The numbers in the brackets denote variations of one standard deviation. Kinesthetic demonstrations were on average longer and had greater variance in length. In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-D, which a human demonstrator cannot emulate consistently.
@@ V Limitations and Future Work
The quantity, quality and diversity of the collected grasps in PvP are partly determined by the viewpoint of the camera. In this work, we empirically select a viewpoint that is adequate for our use-case. It would be interesting to allow the manipulator to explore and find better grasp poses. Furthermore, while tactile sensing is used to improve the robustness of data collection, it could also be used as an extra modality for the policy. We have introduced a framework for data collection and tested it with two real-world tasks. It would be interesting to test PvP on a larger set of tasks. Lastly, PvP is a language-driven data collection method, and it would be natural to integrate it into a high-level language based planner [9].
@@ VI Conclusion
We presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks. PvP leverages recent advancements in grasp planners, tactile sensing, compliant control and the symmetry between picking and placing to autonomously collect placing demonstrations. PvP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by PvP surpasses that of human demonstrations collected by kinesthetic teaching. PvP provides a promising and pragmatic direction for real world robotic placing demonstration data collection with minimal human labour.

### v4

@@ III-A4 Placing by Reversing
An example of a place trajectory is shown in Fig. 1(d). Additionally, we convert the global poses to relative poses in order to use state differences as expert actions:
$\Delta\mathbf{T}_{i}=\bar{\mathbf{T}}_{\bar{M}-i}^{-1}\bar{\mathbf{T}}^{\phantom{-1}}_{\bar{M}-i-1},$ (4)
where $i\in\{0,...,\bar{M}-1\}$ and $\Delta\mathbf{T}_{i}\in\mathrm{SE}(3)$ is a relative change in EE pose. We define the expert action as $\mathbf{a}_{i}=(\Delta\mathbf{T}_{i},a_{\text{gripper},i})$, where $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}-1}$, and store the sequences of observations and actions as an expert trajectory.
We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$. We downsample the images to a resolution of 128 $\times$ 128 and also stack the three previous RGB frames, which results in $\mathbf{o}_{i}\in\mathbb{R}^{128\times 128\times 12}$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}-1\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper (i.e. $\Delta\mathbf{T}_{i}=\boldsymbol{\mathbf{I}}$ and $a_{\text{gripper},i}=0$) for $N$ time steps. We use $N=$ 5, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples
@@
Figure 3: Tactile images from before and after a tactile regrasp (TR). Left: the contact surface area of the plate fills half of the tactile image, indicating a shallow grasp. Right: after a regrasp, the contact surface area has increased, indicating a stable grasp.
Figure 4: Visualization of contact surface region in the tactile image of grasps of various objects: (a) a red metal plate, (b) a red metal mug, (c) a red metal bowl and (d) a green wheat straw plate.
III-B Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first 75% of the poses with noise. We take inspiration from previous works that have investigated the use of noise injection for both the policy observations [38], and actions [39]. We also take inspiration from robust control theory, where injecting isotropic Gaussian noise has been shown to achieve persistent excitation [40], a condition where the training data is informative enough to learn a model that is robust to compounding errors during deployment. We call this variant noise-augmented data collection. In practice, we represent pose commands $\Delta\mathbf{T}$ as a combination of a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a rotation vector $\boldsymbol{\theta}\in\mathbb{R}^{3}$. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We perturb the translation of the pose with
@@ III-B Noise-Augmented Data Collection
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is an isotropic Gaussian noise vector and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation with
@@
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is an isotropic Gaussian noise vector and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta\in\mathbb{R}$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We perturb both the rotation axis and the rotation angle. We find $\sigma_{t}=$ 0.5 cm, $\sigma_{e}=$ 0.5 cm and $\sigma_{\theta}=$ 0.5 to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-C Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}\mid\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by PvP. We use behavioural cloning (BC) [41] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
$\mathcal{L}=\frac{1}{N_{\text{train}}}\sum_{i=1}^{N_{\text{train}}}\sum_{t=0}^{T}-\log{\pi_{\phi}(\mathbf{a}_{t}\mid\mathbf{o}_{t})}.$ (8)
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [42], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}\!\!\mid\!\!\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [43] and [44], by setting the standard deviation of the Gaussian distributions to be an arbitrarily small value. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss.
@@ IV Experiments
(a) Dishrack placement task, which consists of placing multiple plates of varying physical properties in evenly spaced slots.
(b) Table placement task, which consists of placing a bowl on a plate and a cup on a coaster.
@@
We used PvP to collect data and train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Fig. 5. In both cases, the place policy had to account for the weight and shape of each item, the initial grasp of the object, and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we only present our quantitative results for the more challenging dishrack loading task as it better represents the type of task (with added contact constraints from the environment) that PvP was designed for. We study the robustness of PvP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B, and the relative quality of policies trained with data from PvP and from kinesthetic teaching in Section IV-C.
@@ IV-A PvP Data Collection Robustness
We require a robust data collection loop to collect a large number of place demonstrations autonomously with PvP. We studied the effects of CCG and TR, as introduced in Section III, on the robustness of PvP.
We observed two main causes of misplaced objects during data collection, initially discussed in Section III-A2, First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the environment. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the environment. With PvP, we mitigate this failure mode with CCG. The second failure mode involved unstable grasps that lead to a large relative motion between the object and the gripper. With PvP, we mitigate unstable grasps using TR to preemptively detect unstable grasps and to regrasp if needed.
We conducted an ablation study on CCG and TR where we ran self-supervised data collection for a total of 128 episodes and recorded the number of failures. We visualize the results in Fig. 6. The naive approach without CCG and without TR failed a total of 15 times, which roughly translates to an average of nine episodes collected autonomously before a human intervention is needed. The number of failures dropped to three with the addition of CCG, which raises the average to 43 episodes collected before each intervention. Finally, with both CCG and TR, we were able to successfully collect 128 episodes without any human intervention.
@@ IV-B Noise Augmentation Ablation
TABLE I: Success rates for a model trained on a dataset collected with and without noise augmentation and using two different policy action representations. The noise-augmented data collection procedure improves the performance of both deterministic and Gaussian mixture policies.
@@
In Section III-B, we introduced a noise-augmented data collection variant of PvP. We investigated its contribution to the success rate of the policy. We collected two different datasets each containing 128 demonstrations, with and without added noise augmentation. We tested the effect of noise-augmented data on policies trained using two different action representations: a deterministic policy and a Gaussian mixture policy with five modes. We present the results in Table I. For each row, we averaged the success rate over three models trained from scratch with three different seeds. We used 20 or 30 rollouts to calculate the success rate of each model. The numbers in the brackets denote variation of one standard deviation. We also report the total successes summed over the three models. For both action representations, we found that training on noise-augmented data improved the performance of the final policy significantly. The deterministic policy had an improvement of approximately 10% while the mixture policy had a larger improvement of 15%.
We hypothesize that adding noise further accentuates the multimodal nature of the data (i.e., multiple valid actions are available for roughly the same observation) and that a mixture representation for the actions allows the policy to better capture this by not having to average over multiple possible expert actions [45]. We also qualitatively observed that policies trained with noise-augmented datasets were better at re-adjusting the object with respect to the goal.
@@ IV-C Comparison to Kinesthetic Teaching
We compared the performance of policies trained using demonstration data from PvP and from kinesthetic teaching. In particular, we investigated whether the quality of the demonstrations differ. To compare, we trained two separate policies, both sharing the architecture outlined in Section III-C, on two separate datasets: one collected with PvP and another collected with kinesthetic teaching [46]. We used a Gaussian mixture model as the action representation of both policies. For PvP, we used the noise-augmented variant since it was the best-performing model, as shown in Section IV-B.
In Fig. 7, we visualize the success rates of both models trained with varying numbers of demonstrations. As done previously, we averaged the performance over three models with different seeds for each datapoint. We tested each model using eight rollouts. The models trained with data from PvP, collected without any human involvement, significantly outperform the models trained using kinesthetic data by about 20% for almost all numbers of demonstrations. We observed that policies trained with kinesthetic teaching often misplaced the plate (e.g., by not having the right place angle or not resting the plate in a stable manner with respect to the environment contact points) and struggled to open the gripper at the correct times. The latter of the two error modes did not appear in policies trained with PvP.
We hypothesize that the kinesthetic policy performed more poorly than the PvP policy because the overall quality of the human demonstrations was lower than the programmatic or machine-generated demonstrations produced by PvP. Previous work has studied the difficulties and challenges of training policies using human demonstrations when compared to programmatic or machine-generated demonstrations [47, 48, 44]. In particular, human demonstrations can be sub-optimal due to mistakes (e.g., large variations in trajectory length, unwanted movements, and remaining idle for some time to think). The authors of [44] note that the relative average trajectory length is a good proxy for the quality of the dataset. The average number of time steps for all demonstrations was 29.41 (2.22) for the PvP dataset and 41.66 (5.69) for the kinesthetic dataset. The numbers in the brackets denote variation of one standard deviation. Kinesthetic demonstrations were on average longer and had greater variance in length. In addition, PvP benefits from systematic noise-augmentation, as covered in Section III-B, which a human demonstrator cannot emulate consistently.
@@ V Limitations and Future Work
PvP relies on effective grasping (i.e., accurate grasp detections and stable grasps during retrieval). This makes this method susceptible to the typical challenges involved in grasping, namely dealing with noisy point cloud measurements, unreliable object segmentation, and unmodelled object dynamics (e.g. object slip). The quantity, quality, and diversity of the grasp poses found by the planner are partly determined by the viewpoint of the camera. In this work, we empirically selected a viewpoint that is adequate for our use case. It would be interesting to allow the manipulator to explore and find better grasp poses. Furthermore, while tactile sensing is used to improve the robustness of data collection, it could also be used as an extra modality for the policy. We have introduced a framework for data collection and tested it with two real-world tasks. We would like to test PvP on a larger set of tasks for a more exhaustive evaluation. Lastly, PvP is a language-driven data collection method, and it would be natural to integrate it into a high-level language based planner [11].
@@ VI Conclusion
In this paper, we presented PvP, a self-supervised data collection method to collect expert demonstrations for placement tasks. PvP leverages recent advancements in grasp planners, tactile sensing, compliant control, and the symmetry between picking and placing to autonomously collect placing demonstrations. PvP is able to robustly collect a large amount of data consistently without human intervention by making use of compliant manipulator control and regrasping based on tactile sensing. We showed that the quality of the demonstrations produced by PvP surpasses that of human demonstrations collected by kinesthetic teaching. PvP provides a promising and pragmatic direction to collect real world robotic placing demonstration data with minimal human labour.